Avis rents a car for a daily rate of $17.99 plus 18 cents per mile. Hertz charges $18.95 per day plus 16 cents per mile. For what mileage is the cost the same?
Found 2 solutions by ewatrrr, Gersid:
Answer by ewatrrr(24785)


Hi
Let x represent the number of miles in one day:
Question Asks*** For what mileage is the cost the same?
17.99 + .18x = 18.95 + .16x
Solve for x
.02x = .96
x = 48 mi
CHECKING our Answer***
17.99 + .18*48 = 18.95 + .16*48 = 26.63

Answer by Gersid(33)

The charge for Avis is:
A = 17.99 0.18M (M = number of miles)
For Hertz:
H = 18.95+0.16M
Set A = H and solve for M
17.99+0.18M = 18.95+0.16M Subtract 0.16M from both sides.
17.99+0.02M = 18.95 Subtract 17.99 from both sides
0.02M = 0.96 Divide by 0.02
M = 48
The cost would be the same for 48 miles.